Consider a free market for a good with demand equal to Qe 900-10P and supply equal to Q-20P 5. Draw the graph of demand and supply curve. What are the equilibrium price and quantity on the market? a.
Answer #1

Q = 900 - 10P

10P = 900 - Q

P = 90 - 0.1Q (this is the inverse demand function)

When, Q = 0, P = $90

For equilibrium

Demand = supply

900 - 10P = 20P

30P = 900

P = 900 / 30 = $30 (equilibrium price)

Q = 900 - 10(30) = 600 (equilibrium quantity)